oregonduck (oregonduck):

HELP WITH FLVS ALGEBRA 1

oregonduck (oregonduck):

A quadratic equation is shown below: 9x2 - 36x + 36 = 0 Part A: Describe the solution(s) to the equation by just determining the radicand. Show your work. (3 points) Part B: Solve 2x2 - 9x + 7 = 0 using an appropriate method. Show the steps of your work and explain why you chose the method used.(4 points) Part C: Solve 3x2 - 12x + 2 = 0 by using a method different from the one you used in Part B. Show the steps of your work. (3 points)

OpenStudy (vera_ewing):

What exactly do you need help with?

oregonduck (oregonduck):

Part A: Describe the solution(s) to the equation by just determining the radicand. Show your work. (3 points) .....9x² - 36x + 36 = 0 .....9(x² - 4x + 4) = 0 .....9(x-2)(x-2) = 0 .....Radicand = Discriminant = 36² - 4(9)(36) = 0 .....⇒ there is 1 real zero of miltiplicity 2 (in other words, .....there is 1 real zero that repeats itself) .....That zero is x=2. Part B: Solve 2x^2 - 9x + 7 = 0 using an appropriate method. Show the steps of your work and explain why you chose the method used.(4 points) .....2x²-9x+7 = 0..........// Use factoring. Factors of 7 are 1 and 7 .....(2x-7)(x-1) = 0 ..........2x-7 = 0............(x-1)=0 ..........2x = 7................x=1 ..........x = 7/2 .....Your solutions are x=1 and x=7/2 Part C: Solve 3x^2 - 12x + 2 = 0 by using a method different from the one you used in Part B. Show the steps of your work. (3 points) .....3x² - 12x + 2 = 0 ............-(-12)±√((-12)²-4(3)(2)) .....x = ---------------------------------....// Solve by Quadratic formula .....................2(3) ............12±√(144-24) .....x = -------------------- .....................6 ............12±√120 .....x = ------------- .....................6 ............12±2√30 .....x = ------------- .....................6 .....x = 2±(1/3)√30